Q23 WXYZ is a quadrilateral such that WZ = 2XW, ZY = XW, XY = XW + WZ, and XY + ZY = 16. What is
the perimeter of WXYZ?
A) 48
B) 42
C) 36
D) 28

Answer by greenestamps(13200)



WZ = 2XW: let XW = x; then WZ = 2x
ZY = XW: ZY = x
XY = XW + WZ: XY = x+2x = 3x
XY + ZY = 16: 3x+x = 16 --> 4x = 16 --> x = 4

The perimeter is WX+XY+YZ+ZW = x+3x+x+2x = 7x = 28, answer D.